If a sample of size n = 25 is chosen from a normal population, which one of the following is a 95% confidence interval for the population mean when the sample mean is 50 and the sample standard deviation is 6.5? O A 50 2.06(1.3) B 50 2.06(6.5) C 50 2.064(65) D 50 t 2.064(1.3)
Answer #1

Solution :

Given that,

\bar x = 50

s = 6.5

n = 25

Degrees of freedom = df = n - 1 = 25 - 1 = 24

At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,24 = 2.064

standard error = s / \sqrtn = 6.5 / \sqrt25 = 6.5 / 5 = 1.3

The 95% confidence interval for \mu is ,

\bar x +/- t\alpha /2,df * S.E.

50 +/- 2.064(1.3)

Option D) is correct .